OpenStudy (anonymous):

Solve using quadratic formula: x(x-2) - 3(2x-5) = 2 ...Let's just say I'm just as bad with this formula as I am with fractions.

OpenStudy (campbell_st):

well expand and simplify 1st x^2 - 2x - 6x + 15 = 2 so x^2 -8x + 13 = 8 GQF \[(-b \pm \sqrt{b^2 - 4ac})/2a\] a = 1, b = -8 c = 13 ... substitute and simplify. You should get \[(-8 \pm \sqrt{10})/2\]

OpenStudy (campbell_st):

simplifyed equation should be x^2 - 6x + 13 = 0

OpenStudy (anonymous):

The equation is simplified to \[x^2-8x+5=0\] a=1 b=-8 c=5 Quadrstic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (campbell_st):

actuually c = 13... not 8

OpenStudy (anonymous):

Wait... Wouldn't that be: x^2 - 8x + 13 = 0 ?